OpenStudy (anonymous):

If G(x) = 3x + 1, then G-1(1) is

OpenStudy (anonymous):

are you asking for the inverse?

OpenStudy (anonymous):

idk but there are three answers to choose from 4, 1/3 , or 0

OpenStudy (anonymous):

Given that F(x)=x²+2, evaluate F(2). whats the answer for this?

OpenStudy (anonymous):

y = 3x+ 1 3x = y-1 x = (y-1)/3 G-1 = (x-1)/3 G-1 (1) = 0

OpenStudy (anonymous):

so what is it 0 or 1/3?

OpenStudy (anonymous):

0

OpenStudy (anonymous):

0, that was my bad

OpenStudy (anonymous):

thanks ;)

OpenStudy (anonymous):

G-1 (1) = (1-1)/3 = 0/3 = 0

OpenStudy (anonymous):

for your second question you just plug n 2 where you have x f(x)=(2)^2+2 f(x)=6

OpenStudy (anonymous):

Given A = {(1, 3)(-1, 5}(6, 4)}, B = {(2, 0)(4, 6)(-4, 5)(0, 0)} and C = {(1, 1)(0, 2)(0, 3)(0, 4)(-3, 5)}, answer the following multiple choice question: From the list of sets A, B, and C above, choose the set of relations that correctly represents a function. Set A only Sets A and C only Sets A and B only

OpenStudy (anonymous):

I believe it is set A only because a function is defined as being one to one which means when you plug something in you should get a differnet number. So basically you will never have the same x adn y

OpenStudy (anonymous):

Given A = {(1, 3)(-1, 5)(6, 4)}, B = {(2, 0)(4, 6)(-4, 5)(0, 0)} and C = {(1, 1)(0, 2)(0, 3)(0, 4)(-3, 5)}, answer the following multiple choice question: From the list of sets A, B, and C, state the domain and range of set B. Domain of set B: {0, 6, 5} Domain of set B: {2, 4, -4, 0} Set B does not have a domain

OpenStudy (anonymous):

please someone answer that question for me/

OpenStudy (anonymous):

Well I wanna say all real numbers but given the choices I would go with the second choice beacuse domain is where you are looking at the x values. But I am not too sure on this one. Sorry